Solve the following linear equation:
(x+1)/(2x-1)=4/3
Thanks in advance.

Answer by lwsshak3(11628)

Solve the following linear equation:
(x+1)/(2x-1)=4/3
Thanks in advance.
When you have only two terms like the given expression, the best way to start is to cross multiply, that is, use the old rule,"product of the means=the product of the extremes"
x+1)/(2x-1)=4/3
3x+3=8x-4
5x=7
x=7/5
The conventional way to start is to eliminate fractions by multiplying each term by the LCD
The LCD in this equation is 3(2x-1). As you can see, it requires a little more work. However, when you have several or more terms in the equation, this is the method to use.